Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Combinations - Identical vs Non-Identical

Expert replies
by santhoshsram » Thu Jan 12, 2012 9:36 pm
Kurt, a painter, has 9 jars of paint:
4 are yellow
2 are red
rest are brown

Kurt will combine 3 jars of paint into a new container to make a new color, which he will name accordingly to the following conditions:

Brun Y if the paint contains 2 jars of brown paint and no yellow
Brun X if the paint contains 3 jars of brown paint
Jaune X if the paint contains at least 2 jars of yellow
Jaune Y if the paint contains exactly 1 jar of yellow

What is the probability that the new color will be Jaune

a) 5/42
b) 37/42
c) 1/21
d) 4/9
e) 5/9

Source: https://www.beatthegmat.com/difficult-gm ... 8.html#116

I started solving this problem and was struck with a fundamental doubt. If there are n identical objects and if we are to make groups of r objects (r < n), how many groups (or combinations) are possible?

If the objects are not identical, then it is nCr = n! / ((n - r)! x r!). But what if the objects are identical, is it the same?

Take for example, the 4 yellow jars in the question.

Case 1: If they are non-identical i.e. Y1, Y2, Y3, Y4 then the number of ways of picking 3 jars is 4C3.

Case 2 (this is what I'm confused about): If the yellow jars are all identical, Y, Y, Y and Y. Any 3 that I pick are going to be the same YYY and I won't be able to differentiate which 3 these are. So is the number ways of picking 3 jars out of the 4 identical ones in the case 1?

I would really appreciate help from any of the experts.
-- Santhosh S
Join the discussion
Source: — Problem Solving |

by Anurag@Gurome » Thu Jan 12, 2012 10:33 pm
Probability that the new color will be Jaune = 1 - probability that Kurt does not use yellow color
Probability that Kurt does not use yellow color = 5C3/9C3 = 5/42
So, required probability = 1 - 5/42 = [spoiler]37/42[/spoiler]

The correct answer is B.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by GMATGuruNY » Thu Jan 12, 2012 10:33 pm
santhoshsram wrote:Kurt, a painter, has 9 jars of paint:
4 are yellow
2 are red
rest are brown

Kurt will combine 3 jars of paint into a new container to make a new color, which he will name accordingly to the following conditions:

Brun Y if the paint contains 2 jars of brown paint and no yellow
Brun X if the paint contains 3 jars of brown paint
Jaune X if the paint contains at least 2 jars of yellow
Jaune Y if the paint contains exactly 1 jar of yellow

What is the probability that the new color will be Jaune?

a) 5/42
b) 37/42
c) 1/21
d) 4/9
e) 5/9
P(Jaune) = 1 - P(not Jaune).
The new color will not be Jaune if it omits yellow.

P(not Jaune):
P(first jar is not yellow) = 5/9. (9 jars, 5 of them not yellow)
P(second jar is not yellow) = 4/8. (8 jars left, 4 of them not yellow)
P(third jar is not yellow) = 3/7. (7 jars left, 3 of them not yellow)
To determine the probability that all of these events happen together, we multiply the fractions:
5/9 * 4/8 * 3/7 = 5/42.

P(Jaune):
1 - 5/42 = 37/42.

The correct answer is B.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by pemdas » Thu Jan 12, 2012 11:04 pm
i think, Santosh is asking about whether with no difference between the jars of the same color the solution is different. The general assumption here is that the jars of the same color are different/unique. Otherwise you have different approach,

to select 3 jars of the same color --> 3 ways (all brown, red or yellow)
to select 3 jars of different colors --> 1 way (yellow-brown-red in different combination is the same)
to select 2 colors the same and 1 is different --> 6 ways (brown-brown-x, yellow-yellow-x, red-red-x)
:: total 10 ways

to select 'Jaune X if the paint contains at least 2 jars of yellow' --> 3 ways
to select 'Jaune Y if the paint contains exactly 1 jar of yellow' --> 3 ways
:: total 6 ways

Thus, it makes 6/10 or 3/5

the result we obtained is not among the listed choices, and I keep my assumption here that the jars of the same color are different/unique.
Success doesn't come overnight!
Join the discussion